OpenStudy (anonymous):

Is 24:8 in symplest form ?

OpenStudy (eyust707):

is there a number that goes into both?

OpenStudy (eyust707):

hehe

OpenStudy (anonymous):

Wow you help alot :) Thanks

OpenStudy (anonymous):

I think so

OpenStudy (eyust707):

which one?

OpenStudy (anonymous):

is it 2 or 4

OpenStudy (eyust707):

well lets try 4

OpenStudy (anonymous):

KK

OpenStudy (eyust707):

does 4 go into 24?

OpenStudy (kainui):

Try both 2 and 4 and look at it again and see if there are more numbers that you can try out.

OpenStudy (anonymous):

yes 4 can go into 24

OpenStudy (eyust707):

how many times?

OpenStudy (anonymous):

6

OpenStudy (eyust707):

great!

OpenStudy (eyust707):

what about 8 does 4 go into 8?

OpenStudy (kainui):

divide both sides of the ratio by 4 now and tell us your new simplified ratio. Is that completely simplified?

OpenStudy (anonymous):

yep 2 times

OpenStudy (eyust707):

sweet so we get 6:2

OpenStudy (anonymous):

Oh i get it :) Whoop whoop

OpenStudy (eyust707):

butttt is there something you notice...

OpenStudy (anonymous):

iit can be simplyfied

OpenStudy (kainui):

Exactly!

OpenStudy (eyust707):

=D

OpenStudy (eyust707):

is there a number that goes into both?

OpenStudy (eyust707):

what is it and how many times does it go into each?

OpenStudy (anonymous):

2

OpenStudy (eyust707):

yep.. so what do we end up with when we use 2?

OpenStudy (anonymous):

and it gose in once

OpenStudy (eyust707):

2 goes into 2 once how many times does it go into 6?

OpenStudy (anonymous):

3 times

OpenStudy (anonymous):

so is it 2:3

OpenStudy (eyust707):

yayy! so if we have 6:2 what can that be reduced to?

OpenStudy (anonymous):

2:3

OpenStudy (anonymous):

ohhhh 3:2

OpenStudy (eyust707):

ding ding ding

OpenStudy (anonymous):

Yeaaahhhhhh

OpenStudy (sasogeek):

3:1

OpenStudy (eyust707):

ahh yess we made a little mistake

OpenStudy (sasogeek):

remember 2 goes into 2 once...

OpenStudy (eyust707):

how many times does 2 go into 2?

OpenStudy (eyust707):

whats up saso!

OpenStudy (sasogeek):

Note that x:y is the same as x/y ;) 24:8 = 24/8 =3 but every integer is divided by 1... hence 3/1 = 3:1

OpenStudy (sasogeek):

not much eric